### v1

# Snail-Radar: A large-scale diverse dataset for the evaluation of 4D-radar-based SLAM systems

Abstract
4D radars are increasingly favored for odometry and mapping of autonomous systems due to their robustness in harsh weather and dynamic environments. Existing datasets, however, often cover limited areas and are typically captured using a single platform. To address this gap, we present a diverse large-scale dataset specifically designed for 4D radar-based localization and mapping. This dataset was gathered using three different platforms: a handheld device, an e-bike, and an SUV, under a variety of environmental conditions, including clear days, nighttime, and heavy rain. The data collection occurred from September 2023 to February 2024, encompassing diverse settings such as roads in a vegetated campus and tunnels on highways. Each route was traversed multiple times to facilitate place recognition evaluations. The sensor suite included a 3D lidar, 4D radars, stereo cameras, consumer-grade IMUs, and a GNSS/INS system. Sensor data packets were synchronized to GNSS time using a two-step process: a convex hull algorithm was applied to smooth host time jitter, and then odometry and correlation algorithms were used to correct constant time offsets. Extrinsic calibration between sensors was achieved through manual measurements and subsequent nonlinear optimization. The reference motion for the platforms was generated by registering lidar scans to a terrestrial laser scanner (TLS) point cloud map using a lidar inertial odometry (LIO) method in localization mode. Additionally, a data reversion technique was introduced to enable backward LIO processing. We believe this dataset will boost research in radar-based point cloud registration, odometry, mapping, and place recognition.
keywords:
4D radar, odometry and mapping, sensor synchronization, radar extrinsic calibration, data reversion, point cloud registration
1 Introduction
Traditionally, automotive radars are integral to vehicles for obstacle detection and avoidance. In recent years, millimeter-wave radars have also been utilized for odometry and mapping (Harlow et al., 2023) thanks to their robustness to adverse conditions. Two types of radars are commonly used: scanning radars, which rotate to capture 360-view scans (Cen and Newman, 2018), and solid-state single-chip radars, typically with a horizontal field of view of approximately 120 (Cai et al., 2022; Zhuang et al., 2023). Single-chip radars offer higher capture frequencies and are less bulky compared to their scanning counterparts. With the Doppler velocity, a single-chip radar can easily detect moving objects, thus addressing significant challenges encountered in odometry approaches based on cameras (Wang et al., 2021) or lidars (Li et al., 2023). However, conventional automotive radars often have poor vertical resolution, leading to the blending of low and high objects within the same view.
Recent advances in single-chip 4D imaging radars have enabled accurate elevation measurement thanks to improved vertical resolution along with range, azimuth, and Doppler velocity (hence 4D). Utilizing millimeter waves, 4D radars can easily distinguish moving objects and perceive normally in adverse conditions, including fog, rain, and snow. These features make 4D radars appealing for autonomous systems like robots and cars, which must operate reliably in such dynamic and adverse environments.
Despite the surge in research utilizing 4D radars for ego-motion estimation, place recognition, and area mapping, existing public 4D radar datasets, e.g., (Kramer et al., 2022; Li et al., 2022; Choi et al., 2023; Zhang, 2023), often fall short in aspects such as accuracy of reference trajectories, diversity in data collection platforms, geographic scope, and repetition of data acquisition on same routes.
To address these deficiencies, we introduce a curated dataset from a data collection initiative by the SNAIL group at Wuhan University, which began in August 2022. This dataset encompasses a wide array of data sequences collected over a year using diverse platforms including a handheld rig, an e-bike, and an SUV. Though sensor configurations slightly vary across platforms, each sensor rig generally includes a stereo camera, one or two 4D radars, a 3D lidar, one or two IMUs, and a GNSS/INS system. The published sequences were collected in a range of environments from a densely vegetated university campus to highways with tunnels and viaducts, during clear or rainy days, and nighttime.
Considering the applications in simultaneous localization and mapping (SLAM), we provide sequences with reference trajectories generated using data from a terrestrial laser scanner (TLS). Synchronization based on proven techniques and meticulous calibration ensure the reliability of our data. We believe this dataset will significantly aid in evaluating algorithms for odometry, mapping, and place recognition based on 4D radar point clouds. The dataset and accompanying tools are available at our website 11endnote: 1The dataset website https://snail-radar.github.io/.
Our contributions are summarized as follows:
  1. 1.
    We release a large-scale diverse 4D radar dataset, captured at multiple times over selected routes by three different platforms, encompassing diverse environmental conditions including rainy days and nights, campus roads, and highways. Tools for data conversion between a ROS bag and a folder format, along with calibration results, are also provided.
  2. 2.
    Rigorous procedures are proposed and employed to synchronize motion-related messages of all sensors and calibrate the extrinsic parameters between them. The sync procedure starts with the lidar and GNSS sync in hardware, then all sensor times are mapped to GNSS times with lidar times as the bridge, and finally, the constant time offsets between all motion-related message types are estimated. The extrinsic parameters are initialized with manual measurements and refined through a correlation method.
  3. 3.
    A localization pipeline is proposed to generate the reference poses for the provided sequences. The pipeline starts with capturing point clouds of the test area within a campus using a TLS. The stitched point clouds are used as submaps in a LIO method which operates in localization mode. To easily generate reference poses towards the end of a sequence, we also propose a technique to reverse messages for lidar-based odometry.
The structure of the paper is as follows. Section 2 reviews existing radar datasets. The procedure for obtaining 4D radar point clouds is briefly described in Section 3. Details of the dataset, including data collection procedures, sensor setup, and file formats, are provided in Section 4. Ground truth generation, data synchronization, sensor calibration, and known issues are discussed in Sections 5, 6, 7, and 8, respectively. Finally, Section 9 summarizes the paper.
2 Related work
This section reviews existing radar datasets for odometry and mapping with an emphasis on 4D radars. A comprehensive survey of radar datasets targeting various applications including object detection and tracking have been given in (Zhou et al., 2022).
For the three types of radars, spinning radars, classic 3D automotive radars, and the recent 4D automotive radars, a variety of datasets have been published. A detailed comparison of our dataset to the existing datasets is given in Table 1.
With spinning radars, several research groups have released radar localization datasets, e.g., the urban Oxford radar robotcar dataset (Barnes et al., 2020), the labeled RADIATE perception dataset (Sheeny et al., 2021), the urban MulRan place recognition (Kim et al., 2020), and the multi-season Boreas dataset (Burnett et al., 2023).
For the traditional 3D radars, the EU long-term odometry dataset (Yan et al., 2020) employed a Continental ARS308 radar for tracking objects. Also, the NuScenes dataset (Caesar et al., 2020) provides sequences of 20 seconds long captured by five automotive radars. Its accurate poses are obtained by a Monte Carlo localization method relative to the HD map using both lidar and odometry data.
For the recent 4D radars, several datasets have been released. The OdomBeyondVision dataset (Li et al., 2022) was collected by an MAV, a UGV, and a handheld platform in buildings. All three platforms include a FLIR thermal camera and up to three single-chip TI AWR1843 4D radars. The NTU4RadLM dataset (Zhang, 2023) used an Oculii Eagle 4D radar and an iRay thermal camera mounted on a handcart and a car to capture data in clear weather conditions within a campus. The RRxIO dataset (Doer and Trommer, 2021) included a few indoor and outdoor sequences acquired by a thermal camera and a IWR6843 radar mounted on an MAV. The reference trajectories were generated by a motion capture system and a visual inertial SLAM system. The Coloradar dataset (Kramer et al., 2022) was captured by a handheld platform outfitted with a single-chip radar and a cascaded radar. The dataset spans buildings, mine tunnels, and outdoor areas, with reference trajectories generated by a lidar SLAM system. One of its imperfections is that some sequences have wrong Doppler measurements. The USVInland dataset (Cheng et al., 2021) was captured by an unmanned surface vehicles with three TI radars, traveling on inland waterways. The sequences with good GNSS RTK solutions are provided for SLAM evaluations. The MSC RAD4R (Choi et al., 2023) dataset consists of many urban and outskirt sequences captured by a sensor rig mounted on a car. Though the sensor rig includes a GNSS/RTK system and a AHRS system, the RTK solutions often have big closure errors in height and the AHRS system gave wrong headings.
Table 1: Our datasets in comparison to existing radar-based SLAM datasets. Our datasets involves three platforms and diverse scenes, covering small to large geographic scopes.
Dataset Radar Data type Ground truth Platform Weather/Light Scenarios
Oxford Radar
RobotCar
Navtech RA
PGO with
GPS/INS
SUV
fog, night,
rain, snow
urban road
RADIATE Navtech RA GPS/IMU car
fog, night,
rain, snow
urban road, park
MulRan Navtech RA
PGO with
GPS
car clear
urban road,
tunnel, campus
Boreas Navtech RA
GNSS/IMU/
wheel fusion
SUV
night, rain,
snow
urban road
EU Long-term Conti ARS 308 2D PC + D GPS-RTK/IMU car night, snow urban road
ColoRadar Two TI AWR
ADC,
3D PC + D
Vicon or PGO handheld clear
mine, office,
outdoors
MSC-RAD4R Oculii Eagle 3D PC + D GPS-RTK car
night, smoke,
snow
urban road
USVInland TI AWR1843 3D PC + D GPS-RTK/IMU boat clear waterway
Odom Beyond
Vision
TI AWR1843 3D PC + D
MoCap or
LOAM
handheld,
UAV, UGV
smoke indoors
NTU4DRadLM Oculli Eagle 3D PC + D
LVI-SLAM +
PGO
car clear campus
Ours
Oculli Eagle
Conti ARS548
3D PC + D
LIO loc. on
TLS map
handheld,
SUV, UGV
night, rain
campus, highway,
tunnels, overpass
In view of existing datasets, our dataset covers diverse environments (e.g., vegetated campus, tunnels, rain, and nights) using three platforms, featuring repetitive traversals of selected paths and accurate reference solutions.
3 Radar signal processing background
This section reviews the typical data acquisition and processing pipeline for the multi-input-multi-output (MIMO) frequency modulated continuous wave (FMCW) 4D radars, mainly referring to the Texas Instruments (TI) products.
As shown in the top diagram of Fig. 1, the transmitter antennas send out radio frequency chirps with increasing frequency. The receiver antennas detect the waves reflected by objects and downconvert the signals by mixing them with the transmitted carrier wave to obtain signals at the intermediate frequency. These signals are then sampled by an analog-to-digital converter (ADC) to get the complex (in-phase and quadrature components, I/Q) ADC samples. These ADC samples are passed to the digital signal processors (DSPs) for further processing.
The digital signal processing, depicted in the bottom of Fig. 1, includes four fundamental components: range processing, Doppler processing, constant false alarm rate (CFAR) detection, and 2D angle of arrival (AoA) processing. Each component is thoroughly illustrated in the TI mmwave SDK documentation. The mathematical principles underlying this pipeline are explained in (Iovescu and Rao, 2017). Here, we provide a high-level functional description of these components, with advanced options omitted for clarity.
Figure 1: (Top) The schematic of the millimeter wave radar’s mechanism, as in a typical 4D radar by Texas Instruments (TI) (Iovescu and Rao, 2017). PA: power amplifier, BPM: binary phase modulation, LNA: low noise amplifier, IF: intermediate frequency, ADC: analog-to-digital converter, DSP: digital signal processor, AMP: amplifier. (Bottom) The simplified radar point cloud generation pipeline from the ADC samples. CFAR: constant false alarm rate. The math symbols are explained in the main text.
Let’s denote the following:
  • $N_{Dc}$: number of Doppler chirps per frame
  • $N_{Rx}$: number of receiver antennas
  • $N_{Tx}$: number of transmitter antennas
  • $N_{s}$: number of ADC samples per chirp duration
  • $N_{Db}$: number of Doppler bins in the fast Fourier transform (FFT)
  • $N_{R}$: number of range bins in the FFT
The range processing routine takes as input ADC chirp samples of I/Q values, performs 1D range FFT and optional direct current (DC) range calibration during the active frame time, and outputs a radar cube. The ADC data has a size of $N_{Tx}\cdot N_{Dc}\cdot N_{Rx}\cdot N_{s}$. The radar cube has a size of $N_{Tx}\cdot N_{Dc}\cdot N_{Rx}\cdot N_{Rb}$.
The Doppler processing routine takes the preceding radar cube, performs 2D Doppler FFT and energy summation during the inter-frame time, and outputs a detection matrix of size $N_{Rb}\cdot N_{Db}$.
The CFAR routine starts with the detection matrix, performs CFAR detection and peak grouping to output a CFAR detection list of objects in the range-Doppler domain.
The 2D AoA processing routine takes the radar cube and the CFAR detection list, performs 2D Doppler FFT on relevant entries of the detected objects, followed by a 2D angle FFT and peak selection by CFAR to determine the azimuth and elevation for the detected objects, resulting in 3D point clouds with 1D Doppler velocity.
4 Dataset
Our dataset, dubbed snail-radar, is collected using three platforms, a handheld device, an e-bike, and an SUV. The three platforms share almost the same sensor rig as the data collection was carried out sequentially across these platforms. The published dataset includes 44 sequences: 6 handheld sequences, 14 e-bike sequences, and 24 SUV sequences. For each sequence, Table 2 specifies the platform, weather and lighting conditions, and the approximate traveled distance and duration.
Table 2: The 44 sequences of our dataset are repeatedly collected at 8 places with three platforms in diverse weather/lighting conditions. The place shorthands bc=basketball court, sl=starlake, ss=software school, if=info faculty, iaf=info and arts faculty, iaef=info, arts, and engineering faculty, st=starlake tower, 81r=August 1 road. The platforms H=handheld, E=ebike, S=SUV. The weather and lighting condition is clear and daytime unless specified. Different to other sequences of the info and arts faculty, the 20240113/2 sequence’s route is anticlockwise.
Place
Setup
Date/Run
Dist.(m)/
Dur.(sec)
Weather/
lighting
bc H 20230920/1 84/74 light rain, night
H 20230921/2 59/83 mod. rain
E 20231007/4 78/51
E 20231105/6 251/144 mod. rain
E 20231105_aft/2 167/104 light rain
sl H 20230920/2 2182/1839 light rain. night
H 20230921/3 2171/1857 mod. rain
H 20230921/5 2015/1731 mod. rain
E 20231007/2 1997/657
E 20231019/1 1919/460 night
E 20231105/2 2045/524 heavy rain
E 20231105/3 2069/537 heavy rain
E 20231105_aft/4 2019/698 light rain
E 20231109/3 1983/546
ss H 20230921/4 736/622 light rain
E 20231019/2 781/533 night
E 20231105/4 895/395 heavy rain
E 20231105/5 967/400 mod. rain
E 20231105_aft/5 826/534 light rain
E 20231109/4 795/533
if S 20231208/4 2228/515
S 20231213/4 2227/494 light rain. night
S 20231213/5 2225/475 light rain, night
S 20240115/3 2223/525 dusk
S 20240116/5 2228/514 dusk
S 20240116_eve/5 2224/462 night
S 20240123/3 2231/535
iaf S 20231201/2 4620/1042
S 20231201/3 4631/946
S 20231208/5 4616/870
S 20231213/2 4603/938 light rain, night
S 20231213/3 4613/875 light rain, night
S 20240113/2 4610/1005 anticlockwise
S 20240113/3 4613/962
S 20240116_eve/4 4609/868 night
iaef S 20240113/5 7283/1509
S 20240115/2 6641/1374
S 20240116/4 6648/1515
st S 20231208/1 275/147
S 20231213/1 276/126 light rain, night
S 20240113/1 541/214
81r S 20240116/2 8554/1433 light rain
S 20240116_eve/3 8521/1293 night
S 20240123/2 8539/1743
More than half of these sequences were captured under special conditions such as rain, dusk, and night. The dataset encompasses various scenes including campus environments, highways, and tunnels. Sample data from camera, lidar, and 4D radar are provided in Table. 3. For place recognition purposes, data for each route were captured at least three different times. Additionally, each sequence was recorded with roughly the same start and end poses.
ZED2i Left Hesai XT32 Oculii Eagle ARS548
D [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
R [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
N [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
H [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
T [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3: Samples of four sensors in different scenarios. D: Day, R: Rain, N: Night, H: Highway, T: Tunnel. Rain causes many speckles in the lidar point cloud. The enhanced point cloud of Oculii Eagle formed by accumulating multiple frames has more points than a single frame of ARS548. The accumulation also leads to some points behind the platform. The tunnel data’s perceptual aliasing causes the lidar(-inertial) odometry to give wrong velocity disagreeing with those from the GNSS/INS system and the two radars.
There are a total of eight routes, as illustrated in Fig. 2: the basketball court, the starlake, the software school, the starlake tower, the info faculty, the info and arts faculty, the info arts and engineering faculty, and the August 1 road. The basketball court is a small flat area surrounded by some buildings. The starlake route traverses the campus, featuring dense vegetation. The software school route passes through urban canyons formed by tall buildings. the starlake tower route encircles the high-rise starlake tower. Since the Wuhan University main campus consists of three segments, the info faculty, the arts and sciences faculty, and the engineering faculty, we collected data on three incremental routes spanning these segments, featuring urban roads in the campus. The August 1 road route is a highway including a long tunnel.
Figure 2: The routes used for data collection in our dataset. The top plot shows four routes: the basketball court (yellow), the starlake (pink), the software school (brown), and the starlake tower (blue). The bottom plot shows the other four routes: the info arts and engineering faculty (pink), the info and arts faculty (yellow), the August 1 road (blue), and the info faculty (brown).
4.1 Sensor setups
We assembled the sensors on a aluminum frame designed with CAD, ensuring it was waterproof. The sensor setups for three platforms are drawn in Fig. 3, with sensors for data collection listed in Table 4. The handheld rig consists of a Hesai Pandar XT32 lidar, an Oculii Eagle radar, a ZED2i stereo camera, and a Bynav X36D GNSS/INS system with a single antenna. The rig mounted on the ebike or the SUV includes a Hesai Pandar XT32 lidar, a Continental ARS548 radar, an XSens MTi3DK IMU, an Oculii Eagle radar, a ZED2i stereo camera, and a Bynav X36D GNSS/INS system with dual antennas. Specifically, sequences captured before Nov 1 2023 do not include data from the ARS548 radar and MTi-3DK IMU. These sequences encompass all handheld sequences and some ebike sequences.
Table 4: The descriptions of sensors in our dataset. The accelerometer noise and random walk are in units of $m/s^{2}/\sqrt{Hz}$ and $m/s^{3}/\sqrt{Hz}$. The gyroscope noise and random walk are in units of $rad/s/\sqrt{Hz}$ and $rad/s^{2}/\sqrt{Hz}$. FOV: field of view, BS: bias stability.
Type Sensor Rate Characteristics
4D Radar
Conti
ARS548
15
max range 300 m
Doppler [-400, 200] m/s
HFOV 120 VFOV 40
range accuracy 0.3 m
az. accuracy 0.2
elev. accuracy 0.1
#point/frame $\sim$150
Oculii
Eagle
15
max range 400 m
Doppler $\pm$86.8 m/s
HFOV 113 VFOV 45
range accuracy 0.86 m
az. accuracy 0.44
elev. accuracy 0.175
#point/frame $\sim$7500
GNSS/
INS
Bynav
X36D
100
real-time GNSS/INS
H 0.235 m RMS
V 0.14m RMS
in 10 s outage
IMU
ZED2i
IMU
400
accel. noise 1.6×$\times 10^{-3}$,
random walk 2.5×$\times 10^{-4}$;
gyro. noise 1.2×$\times 10^{-4}$,
random walk 3.4×$\times 10^{-5}$
MTi3DK 100
roll/pitch 0.5 RMS
yaw 2 RMS
accel. noise 7×$\times 10^{-4}$,
BS 4×$\times 10^{-4}m/s^{2}$;
gyro. noise 5.236×$\times 10^{-5}$,
BS 6/hr
X36D
IMU
100
accel. noise 5.833×$\times 10^{-4}$,
BS 1.5×$\times 10^{-4}m/s^{2}$;
gyro noise 2.91×$\times 10^{-5}$,
BS 1.8/hr
Lidar
Hesai
Pandar
XT32
10
max range 120 m
range accuracy $\pm$2 cm
FOV ${}^{\circ}\times$31
#point/frame $\sim$90000
Stereo
camera
ZED2i 20
1280$\times$720, grayscale,
HFOV $110^{\circ}$, VFOV $70^{\circ}$
A ThinkPad P53 laptop running Ubuntu 20.04 with a 1TB solid-state drive is used for real-time data preprocessing and recording. This laptop supports GPU processing and is compatible with all ROS drivers. It connects to the internet via a WiFi hotspot hosted by a mobile smartphone with 5G networking service. The internet connection is primarily used for updating the computer time via NTP (Network Time Protocol) and receiving RTCM (Radio Technical Commission for Maritime Services) messages used in RTK GNSS positioning, broadcast by the Qianxun FindCM service.
The ARS548, Oculii, and XT32 are connected to the laptop by Ethernet cables, while the ZED2i and MTi3DK are connected via USB Type-C connectors. The X36D connects to the computer through both a USB A connector and an RJ45 connector. The USB A connection (at COM1) is for receiving control commands, and the RJ45 connection is for sending INS solutions and raw data to the computer, as well as receiving RTCM messages relayed by the computer. The X36D’s PPS signals and GPGGA messages at COM1 are directly fed to the Hesai XT32 through a wired connection to its GPS port, ensuring that the lidar messages are synchronized to the GNSS time.
For the XT32, Eagle, MTi3DK, and ZED2i, data is captured as-is by their official ROS drivers. The ARS548 and X36D UDP (User Datagram Protocol) packets are recorded on-site using the tshark utility.
During daytime data collection, the ZED2i’s auto-exposure is enabled, usually adjusting the exposure time to less than 5 ms, often around 2 ms. At dusk or night, auto-exposure is disabled, and the exposure time is locked at 5 ms to reduce motion blur.
All data collections start from an open area. Every time the laptop and all sensors are powered up, we first ensure that the GNSS solution is fixed, then move the platform around to perform INS alignment until the INS solution converges.
(a)                                                                                (b)                                                                                               (c)
Refer to caption Refer to caption
(d)                                                                                                                        (e)
Figure 3: The sensor rig on three platforms and the coordinate frames of the sensors. The Thinkpad P53 laptop and the bundle of cables are concealed in a waterproof bag.
4.2 File formats
For each sequence, a monolithic rosbag as well as a folder containing separate messages as individual files is provided on OneDrive. The message types in a rosbag and their respective folders are listed in Table 5. In these folders, the point clouds from the Pandar XT32 lidar, the ARS548 radar, and the Eagle radar are saved as pcd files for easy visualization. The compressed images of the ZED2i stereo camera are saved as jpg images. The other topics, e.g., GNSS/INS solutions, and IMU data are saved in txt files.
The ARS548 point clouds are decoded from the detection lists in UDP packets. The X36D messages are also decoded from the UDP packets following the Bynav protocol 22endnote: 2Bynav data communication interface protocol UG016 https://www.bynav.com/cn/resource/support?cat=6.
Table 5: The ROS1 topics and message types in the published rosbags, along with the corresponding folders in the published zip files.
Sensor Topic ROS message Folder Format
Conti ARS548 /ars548 PointCloud2 ars548/points pcd
Hesai XT32 /hesai/pandar PointCloud2 xt32 pcd
MTi3DK IMU /mti3dk/imu Imu mti3dk txt
Oculii Eagle /radar_enhanced_pcl2 PointCloud2 eagleg7/enhanced pcd
/radar_pcl2 PointCloud2 eagleg7/pcl pcd
/radar_trk PointCloud eagleg7/trk bin
Bynav X36D /x36d/gnss NavSatFix x36d txt
/x36d/gnss_ins NavSatFix x36d txt
/x36d/imu_raw Imu x36d txt
ZED2i /zed2i/zed_node/imu/data Imu zed2i txt
/zed2i/zed_node/left_raw/
image_raw_gray/compressed
CompressedImage zed2i/left jpg
/zed2i/zed_node/right_raw/
image_raw_gray/compressed
CompressedImage zed2i/right jpg
/zed2i/zed_node/odom Odometry zed2i txt
For each sequence, we provide reference trajectories of the lidar XT32 frame in the TLS map frame at 10 Hz. However, for large SUV sequences, we only provide the reference poses during the starting subsequence and end subsequence which are covered by the TLS point clouds. To be complete, we provide all the TLS point clouds and their refined poses.
For all sequences, we also provide the real-time kinematic (RTK) GNSS/INS solutions, i.e., the latitude, longitude, and ellipsoid height wrt the WGS84 (EPSG4326) reference frame, the used GNSS service, the solution status, and diagonal position covariance entries extracted from the Bynav messages INSPVAXA and RAWIMUSA. The coordinates are also converted to the UTM50N zone (EPSG 32650), since Wuhan is located at the northwest corner of UTM50R zone. The INS solution is sometimes unreliable in the vertical direction, having spikes up to a few meters.
The sensor calibration data are provided in scripts 33endnote: 3Dataset tools https://github.com/snail-radar/dataset_tools.
5 Ground truth
The GNSS RTK/INS solutions from the Bynav system often exhibit height jumps of up to a few meters (Toth and Jóźków, 2016). Therefore, we generate the reference poses using precise TLS point clouds, similar to (Ramezani et al., 2020). But instead of frame-wise ICP, we sequentially align the undistorted lidar frames to the TLS map using a LIO method.
A total of 93 TLS scans were captured by a Leica RTC360 scanner on a sunny day, covering the starlake and the starlake tower routes. In capturing these scans, we ensured sufficient overlap between consecutive scans, with a mean distance of 28.2 m between consecutive scans. These scans were first processed using the Cyclone Register 360 program, with manually identified additional loop constraints between some scans. The TLS scan registration results were further refined by point-to-plane ICP in Open3D and regularized by SE(3) pose graph optimization with uniform weights while considering the two loops within these scans. The pairwise registration results were inspected by three individuals over approximately 10 rounds to ensure proper pairwise matching. The position accuracy of the TLS pairwise registration is expected to be within 5 cm, as implied by the quality check in Fig. 4. The final TLS map was created by stitching together these 93 scans.
Next, for each sequence entirely covered by the TLS map, we localized the lidar scans relative to the TLS map in two steps: first getting the initial pose, and then running LIO in localization mode. To get the initial pose, we selected a lidar scan at the stationary beginning, and registered the scan to a nearby TLS scan in the CloudCompare program by manually picking five pairs of correspondences. The labor is acceptable as our routes used fewer than five unique starting poses separated by a notable distance. Every initial pose was further refined with the classic point-to-plane ICP in Open3D and manually checked in terms of the number of matches and inlier RMSE (root mean squared error). With these initial poses, we run an extended FAST-LIO2 method (Xu et al., 2022) in localization mode. This method loads the constant TLS map and predicts the pose with IMU data, undistorts the lidar scan, and iteratively refines the pose with matches between scan points and map surfels. The LIO method operates in localization mode as the incremental submap building is replaced by loading segments of the TLS map in a sliding manner. Also, to remove the real-time operation constraint and execution randomness, the method was converted to run sequentially in offline mode. The IMU data used in the localization method is from the MTi3DK or the X36D when the MTi3DK is unavailable.
For large-scale sequences, since the TLS map only covers the beginning and end parts of the sequence, we generate the reference trajectory only for the beginning and end subsequences within the TLS coverage, as done in TUM-VI (Schubert et al., 2018). We select data from the first four minutes as the start subsequence and data from the last four minutes as the end subsequence. The start subsequence is processed by the LIO in localization mode which is terminated when the platform’s minimum distance to the TLS trajectories exceeds 8 m. The TLS trajectories are simply the reference poses of multiple medium-scale sequences which cover the entire TLS area.
The end subsequence is processed analogously to the start subsequence. However, the end subsequence is reversed to enable processing from the end, as we only know the initial pose at the end. The reverse sequence is created with the requirement that the platform poses remain unchanged in space despite being reversed in time. Specifically, we first find the maximum time of the sequence $t_{max}$, then modify every message’s stamp and every lidar point’s stamp to $t^{\prime}=2t_{max}-t$ from its original time $t$. The accelerometer data remain intact $\mathbf{a}_{m}^{\prime}\leftarrow\mathbf{a}_{m}$, but the gyro data are inverted in sign $\boldsymbol{\omega}_{m}^{\prime}\leftarrow-\boldsymbol{\omega}_{m}$. After processing with LIO in localization mode, we remove the last 8 seconds of the resulting poses for both subsequences. The end subsequence poses’ times are then restored using $t_{max}$. Finally, the start poses and the end poses are combined to form the reference poses for the large scale sequence.
The backward processing also offers a means to verify the reference accuracy by comparing the forward and backward results. For the 20231105/6 sequence, we plot the positions of the lidar relative to the TLS map, $\mathbf{p}_{WL}$ obtained by the forward and backward processing, and their differences $\delta\mathbf{p}_{WL}$ in Fig. 4. The root mean square of $\delta\mathbf{p}_{WL}$ is 9.5 cm, indicating that the reference positions are mostly accurate within a decimeter.
Figure 4: For the 20231105/6 sequence, positions of lidar frames relative to the TLS map obtained through forward and backward processing using the LIO method in localization mode.
6 Synchronization
Since it is challenging to sync all sensors in hardware, we propose a scheme to ensure that their messages are stamped by the same virtual clock. The scheme first uses the lidar data as a bridge to map all sensor message timestamps to the GNSS time. Then, the constant time offset between message topics is estimated by odometry and correlation algorithms.
[…]
After mapping to the GNSS time, a small constant time offset may persist between two streams of messages, e.g., the Zed2i IMU and the Zed2i images. due to transmission delays. This constant time offset between two motion-related message topics can be found using the correlation method. Since all messages in our dataset are motion-related except the camera information messages, we estimate the time offset of these message streams relative to the X36D GNSS/INS trajectory.
Specifically, for the XT32 lidar, we first process the lidar scans with the lidar odometry (LO) method, KISS-ICP (Vizzo et al., 2023), which gives 3D lidar poses at the beginning of each scan. Next, the angular rates in the lidar frame {L} and the INS frame {I} are computed from the INS poses and LO poses by central difference, obtaining $\boldsymbol{\omega}_{WL}^{L}$ and $\boldsymbol{\omega}_{WI}^{I}$. Finally, the time delay of the lidar data is computed by the Correlation Algorithm 1, given the angular rates and the nominal $\mathbf{R}_{IL}$.
For a 4D radar, e.g., the ARS548 or the Oculii Eagle, we first process the radar scans with the GNC method (Yang et al., 2020) to estimate its ego velocity, $\mathbf{v}_{WR}^{R}$. Next, the ego velocity $\mathbf{v}_{WR}^{I}$ is computed from the INS poses with the nominal lever arm $\mathbf{p}_{IR}^{I}$ by central difference. Finally, the time delay of the radar data is estimated using the correlation algorithm, given the ego velocities and the nominal $\mathbf{R}_{IR}$.
For the IMUs, e.g., the ZED2i IMU, which directly provides the angular velocity, $\boldsymbol{\omega}_{WU}^{U}$ in its frame {U}, we first compute the angular rates in the INS frame, $\boldsymbol{\omega}_{WI}^{I}$, by central difference the INS poses, and then run the correlation algorithm given the angular rates and the nominal $\mathbf{R}_{IU}$.
The correlation algorithm 1 finds the time offset $t_{d}$ and relative rotation $\mathbf{R}_{AB}$ between two stamped sequences $\{\mathbf{v}_{i}^{A}\}$ at $\{s_{i},i=1,\cdots,N_{a}\}$ and $\{\mathbf{v}^{B}\}$ at $\{t_{j},j=1,\cdots,N_{b}\}$, e.g., angular rates or ego velocities. Its idea to find the maximum cross correlation is straightforward, but there are a few parameters worth noting. The default smoothing window sizes $M_{a}$ and $M_{b}$ are 1, meaning no smoothing is applied. When the two sequences have very different rates, e.g., 100 vs 10, a smoothing window of size e.g. 5, is necessary for the higher rate sequence. This smoothing causes signal delay which should be accounted for when computing the time offset $t_{d}$. The time tolerance for associating the time-corrected data should typically be smaller than a fraction of the resampling interval, e.g. $\delta t/2$. The truncated least squares method in each iteration first solves for the rotation perturbation $\delta\boldsymbol{\theta}$ and secondly update $\mathbf{R}_{AB}$ by $\mathbf{R}_{AB}\leftarrow\mathrm{Exp}{(\lfloor\delta\boldsymbol{\theta}\rfloor_{\times})}\mathbf{R}_{AB}$. $\delta\boldsymbol{\theta}$ is obtained by solving the linear system stacked from the basic residual equation,
$\lfloor\mathbf{R}_{AB}\mathbf{v}^{B}\rfloor_{\times}\delta\boldsymbol{\theta}\doteq(\mathbf{R}_{AB}\mathbf{v}^{B})\times\delta\boldsymbol{\theta}=\mathbf{v}^{A}-\mathbf{R}_{AB}\mathbf{v}^{B},$ (1)
while discounting those with large residuals.
Algorithm 1 Correlation Algorithm for 3D Motion Data
1:Two sequences of timestamped velocity data $\{\mathbf{v}_{i}^{A}\}$ at $\{s_{i},i=1,\cdots,N_{a}\}$ and $\{\mathbf{v}^{B}\}$ at $\{t_{j},j=1,\cdots,N_{b}\}$; rough guess of the rotation between the two global frames, $\mathbf{R}_{AB}$.
2:Estimated time offset between two sequences $t_{d}$ and refined $\mathbf{R}_{AB}$.
3:Keep only the overlapping part (including a buffer at both ends) of the two sequences.
4:Smooth the data by a sliding window of Gaussian weights to remove the high frequency components and improve correlation. The window sizes are $M_{a}$ and $M_{b}$ for the two sequences.
5:Determine the sampling interval $\delta t$ as the smaller mean interval of the two sequences, and allocate the sample times for both sequences, $\{s_{i}^{\prime},i=1,\cdots,N_{a}^{\prime}\}$, and $\{t_{j}^{\prime},j=1,\cdots,N_{b}^{\prime}\}$.
6:Cubic fit the two sequences of data and get samples at the proposed sample times.
7:Compute norms of the two sequences of data and cross-correlate the two norm sequences.
8:Locate the peak correlation step, and refine its position $d$ by quadratic fitting.
9:Compute the time offset $t_{d}=d\delta t+s_{1}^{\prime}-t_{1}^{\prime}+(M_{b}-M_{a})\delta t/2$.
10:Correct the timestamps of $\{\mathbf{v}_{j}^{B}\}$ by adding $t_{d}$.
11:Associate $\{\mathbf{v}_{i}^{A}\}$ and $\{\mathbf{v}_{j}^{B}\}$ in time with a tolerance $\tau$.
12:Refine $\mathbf{R}_{AB}$ by the truncated least squares method.
Though the correlation algorithm is robust and globally optimal, it can still fail in certain cases, e.g., when multiple correlation peaks occur. Failed cases are spotted by first checking whether the time offsets along 4 axes (norm, x, y, and z) are consistent and whether the resulting relative rotation is close to the nominal value, and secondly visually checking the time alignment plots. These failures, about 10 out of 300 correlation attempts, were manually corrected case-by-case by tweaking the smoothing window size and/or masking the first peak.
[…]
Overall, our sync precision is expected to be better than 10 ms, considering the following facts. 1. The GNSS/INS solution has a rate of 100 Hz, and the correlation algorithm uses quadratic fitting to refine the maximum correlation position. 2. The reported $\sigma$ of Swift-VIO toward the end of a sequence is always less than 1 ms. 3. The exposure time of ZED2i is adjusted to be no more than 5 ms.
The sequences in the published dataset already have the message timestamps compensated by the estimated time offsets to the GNSS time. Specifically, the stamps of all motion-related messages are corrected by the corresponding time offsets. For points in every XT32 lidar frame, their timestamps are also corrected by the lidar time offset.
7 Sensor rig calibration
For the dataset, we define the body frame to have the same origin as the XT32 lidar, but with an orientation such that its x-axis points forward, y-axis points left, and z-axis points up. Thus, the body frame relates to the XT32 frame by a known rotation. The extrinsic parameters of all sensors are provided relative to the body frame.
Initial relative positions between sensors are measured manually, and initial relative orientations are obtained from the CAD drawing 3. These parameters are then refined through several methods.
The extrinsic parameters of the ZED2i IMU or the MTi3DK relative to the XT32 lidar are refined using the lidar IMU calibration tool (Zhu et al., 2022). For our specific ZED2i, there is an open data sheet providing the camera intrinsic parameters, IMU noise parameters, and camera extrinsic parameters, downloadable from its website. This allows us to relate these ZED2i sensors to the body frame by concatenating these extrinsic transforms.
For the Bynav system, we obtain the refined relative rotation between the XT32 lidar and the Bynav IMU by the correlation algorithm 1.
The relative orientation of the Oculii Eagle and ARS548 radars to the X36D IMU is also refined using the correlation algorithm. For the ARS548 radar mounted on top of the rig, it slightly tilted backward for the few sequences captured on Nov 5 2023. Therefore, we provide its relative orientation for every sequence on that day.
8 Known issues
We have identified several issues in the dataset.
  • To save storage space, the resolution of the ZED2i stereo images has been reduced to 640$\times$360 in resolution since Dec 8 2023.
  • In some sequences, the ZED2i images and IMU data have lower rates than the nominal values, likely due to low power voltage.
9 Conclusions
This paper presents a large-scale 4D radar dataset designed for localization and mapping applications based on multi-sensor fusion. The dataset was collected using three platforms under diverse environmental conditions, including rainy days, nights, campus roads, and tunnels. A total of 44 sequences were repeatedly collected over eight routes.
Reference poses (at the start and end for large-scale sequences) are generated using a LIO method that sequentially localizes the undistorted Hesai lidar frames to the TLS map. A data reversion technique enables backward LIO in time. Sensor data are synchronized using a two-step scheme which first uses the hardware-synced lidar data to map the message host times to GNSS times to remove jitter and long-term drift, and secondly uses odometry and correlation algorithms to estimate the constant time offsets between sensor data and the GNSS/INS solution. The extrinsic parameters between all sensors are refined using several methods and are provided with the dataset. We hope our dataset will facilitate research into the newly advented 4D radars, particularly in odometry, mapping, and place recognition.
{acks}
We acknowledge the assistance of Xiaochen Wang and Hanwen Qi in collecting the TLS data with the RTC360, lent by Professor Xinlian Liang. We are also grateful to Shangjia Liu and Qi Liu for their help with data collection.